what is the number in the series
5,25,61,113,181
plss help fast
pls dont use algbra to explain
thanks

Answer by venugopalramana(3286)

what is the number in the series
5,25,61,113,181
plss help fast
pls dont use algbra to explain
T1=5..T2=25..T3=61..T4=113...T5=181
T2-T1=25-5 = 20=D1=20
T3-T2=61-25 = 36=D2=20+16
T4-T3=113-61= 52=D3=20+2*16
T5-T4=181-113=68=D4=20+3*16
HENCE T6 CAN BE OBTAINED BY USING
T6-T5 =T6-181 = D5 = 20+4*16=84
T6=181+84=265